Kumar is three times as old as his sister. Three years from now, he will be twice as old as his sister.Find their present ages.
Answer by satyareddy22(84)

kumar=k,sister=s
k=3s,
k+3=2(s+3)
3s+3=2(s+3)=2s+6
3s-2s=6-3
s=3yrs
k=3s=3*3=9yrs